OpenStudy (anonymous):

Please help. The volume of a sphere is 4200 ft^3. What is its surface area to the nearest tenth?

OpenStudy (anonymous):

do you know what the formulae for the volume of a sphere is?

OpenStudy (anonymous):

V=4/3(pi)(r^3)

OpenStudy (anonymous):

pefect so you know what the volume is so can you rearange this equation to make r the subject ? Ie r = something

OpenStudy (anonymous):

are you having problems can i help?

OpenStudy (anonymous):

I can't really get it, I can get up to the point that I have to do cubed square roots... I'm not sure how to do that.

OpenStudy (solomonzelman):

@wolfgirl, you just need to find the radius basically.

OpenStudy (solomonzelman):

Sphere Surface Area = 4 • π • r² = π • d² Sphere Volume = 4/3 • π • r³ = ( π •d³)/6 Can you do it now?

OpenStudy (anonymous):

\[4\div 3 \times \Pi \times r ^{3} =4200\]

OpenStudy (anonymous):

\[r ^{3}=4200 \div (4/3 \times \Pi)\] \[r= \sqrt[3]{(3150 \div \Pi}\] \[r= \sqrt[3]{1002.676141}\] r= 10.00891253

OpenStudy (solomonzelman):

WHy are you solving it for her, I wanted her to think!

OpenStudy (solomonzelman):

it is an openstudy not open answers....Please!

OpenStudy (anonymous):

I didn't solve it completely, because she still needs to calculate the surface area. Besides that, she said what she was having difficulties with, which was getting it to the point of solving r^(1/3). If she expressed what she was having a problem with means that she attempted to solve it. If she doesn't know how to do it, how is she supposed to do it? Now, she knows the process and will be able to solve a similar problem by herself.

OpenStudy (anonymous):

Could someone please just tell me how to find a cubed root?